# Transfer matrix approach to quantum systems subject to certain Lindblad evolution

### v1

@@ Abstract
Solving for time evolution of a many particle system whose dynamics is governed by Lindblad equation is hard. We extend the use of transfer matrix approach to a class of Linblad equations that admit a closed hierarchy of two point correlators. An example that we treat is the XX spin chain, i.e., free fermions, subject to the local on-site dephasing, but can be extended to other Hermitian dissipators, e.g., non-local dephasing. We find a simple expression of the Green’s function in the Laplace domain. The method can be used to get analytical results in the thermodynamic limit, for instance, to get the evolution of magnetization density and to explicitly see the cross over between the ballistic and diffusive behavior, or to show that the correlations between operators at distance $l$ decay with time as $1/t^{\lceil l/2\rceil+1/2}$. It also provides a fast numerical method to determine the evolution of the density with a complexity scaling with the system size more favorably than in previous methods, easily allowing one to study systems with $\sim 10^{6}$ spins.
@@ I Introduction
The transfer matrix formalism has been applied to solve wide variety of problems in physics. A couple of notable examples include propagation of electromagnetic, acoustic waves Yeh (2012); Pendry and MacKinnon (1992); Pendry et al. (1992) or waves of quantum particles such as electrons across any scatterers Mello and Kumar (2004); Beenakker (1997); Kirkman and Pendry (1984); Ando (1989); Pendry (1982, 1984); Lee and Joannopoulos (1981a, b) and computing partition functions Izergin et al. (1992); Schultz et al. (1964). This method applies when the solution of the problem can be built iteratively using a using products of matrices. Therefore, it naturally becomes useful while dealing with problems requiring solutions of linear equations having finitely many non-zero off-diagonals, e.g. tri-diagonal or block tri-diagonal matrices Akaike (1973); Molinari (1997); Reuter and Hill (2012) which describe the Hamiltonian of non-interacting single particle systems. Some simple examples of such systems include the short-ranged quantum tight-binding models or classical harmonic systems. For such systems the transfer matrix formalism allows one to obtain the elements of the single particle Green’s function for the system Lee and Joannopoulos (1981b); Reuter et al. (2011), which later on determines properties such as the two-point correlation functions, current, conductance etc Dhar and Sen (2006); Dhar and Roy (2006).
@@
Another class of models where such problems occur are quantum mechanical models in presence dissipation or dephasing due to certain type of Hermitian Lindblad operators Žnidarič (2010); Žunkovič (2014); Barthel and Zhang (2022); Wang et al. (2024); Žnidarič (2011, 2024); Eisler (2011); Temme et al. (2012). In these cases, the dynamics of two point correlation matrix is given by a linear differential equation where the matrix governing the dynamics has finitely many non-zero off-diagonals. One should therefore be able to apply the transfer matrix method in a manner similar to tight-binding lattices to obtain solutions for such equations, and this is the approach taken in this paper. The transfer matrix approach avoids the calculation of the entire spectral decomposition of the matrix governing the dynamics of the correlators and instead expresses the solution as a power of a simple transfer matrix of some small dimension. The latter allows one to take the thermodynamic limit with ease. In this paper, we demonstrate the transfer matrix approach in the simple case of XX spin chain with $L$ spins in presence of local dephasing and show that this approach leads to a simple solution for the two-point correlation matrix. We express the solution for the correlators in terms of $L^{th}$ power of a $2\times 2$ transfer matrix, and therefore the thermodynamic limit just requires determining its largest eigenvalue and eigenvector. Once the correlation matrix is known, its elements determine physical observables of general interest. For example, the diagonal elements of the correlation matrix directly gives the magnetization density and the imaginary parts of the first off-diagonal elements give the magnetization current between the different bonds in the chain. Using the transfer matrix approach, we express the Green’s function for the magnetization density as a rather simple function in the Laplace domain whose analytic structure determines type of the behavior we observe. At long times diffusion arises due to contribution from real poles of the Green’s function while at short times the ballistic behavior comes from the contribution of its branch cut.
@@
XX spin chain with dephasing has been extensively studied in the literature, e.g. Refs. Žnidarič, 2010; Medvedyeva et al., 2016; Teretenkov and Lychkovskiy, 2024; Silva et al., 2023; Cao et al., 2019; Turkeshi and Schiró, 2021; Žnidarič, 2015; Carollo et al., 2017; Bernard and Jin, 2019; Haga et al., 2023, and therefore serves as an ideal simple case that one can consider, and where one can study e.g. a crossover from ballistic magnetization transport at short times to diffusive behavior at long times. One can think of it as the simplest model, that is on top of it solvable, displaying such a transition. In the continuum limit, a qualitative understanding of the crossover can be described by considering two conserved quantities, say energy and particle density, that are coupled through a scattering. In our spin language we will have the magnetization density $m(x)$ and the magnetization current $j(x)$, with the simplest possible two coupled continuity equations being,
@@
which is nothing but the so called telegrapher’s equation, commonly used in signal analysis for transmission and propagation of electrical signals Metaxas et al. (1983). For small times the diffusive time derivative can be neglected and the equation will behave as a wave equation with velocity $v^{2}$, while for long times the second time derivative is negligible and it goes into a diffusion equation with diffusion constant $D=v^{2}/(4\gamma)$ (see e.g. Ref. Kac, 1974 for a nice way to solve telegrapher’s equation). As we shall see, the equations for the XX chain with dephasing are quite similar, and one can view it as a lattice version of the telegrapher’s equation with a tight-binding dispersion relation, i.e., we have particles having different velocities. What we will get out of our analysis is a behavior that continuously transitions with time from ballistic to diffusive transport, or, in mathematical terms in the continuum limit from a hyperbolic to a parabolic partial differential equation.
This paper is structured as follows: In sec. II, we introduce the model and the matrix equation for the two point correlators that we want to solve. In the next section, sec. III, we adopt the transfer matrix approach to obtain the solution for the correlation matrix at time $t$ for any finite size $L$. In the following section, sec. IV we look at the solution in the thermodynamic limit and discuss the asymptotic of the behavior of diagonal and the off-diagonal elements of the correlation matrix. We also present numerical results for the transferred magnetization and its logarithmic derivative and discuss the crossover from the ballistic to diffusive behavior. We conclude in section  V.
@@ II XX-Model with Dephasing
In this section, we introduce the model and setup the equations for the two point correlators that we wish to solve. We consider an $XX$ spin chain with periodic boundary conditions. The Hamiltonian is given by,
@@
Solving the master equation directly requires solving $14^{L}-$ 1 coupled differential equations, and is therefore hopeless for large system sizes. However, certain Eisler (2011); Žunkovič (2014); Barthel and Zhang (2022) Hermitian Lindblad operators with quadratic Hamiltonians allow reducing the exponential complexity in system size to a polynomial complexity. This is achieved by decoupling blocks of observables based on the number of Fermionic operators they contain resulting in a hierarchy of equations. In this hierarchy, lower point correlators serve as source terms for the equations of higher point correlators. For example, the block of two point correlators decouples from the higher point correlators and the two point correlators satisfy a closed set of linear equations. These then serve as source terms for third order correlators and so on.
@@
$A_{j}^{(r)}(t)$ $=\sigma_{1}^{j}Z_{j+1}^{(r-2)}\sigma_{1}^{j+r}+\sigma_{2}^{j}Z_{j+1}^{(r-2)}\sigma_{2}^{j+r},$ (8)
$B_{j}^{(r)}(t)$ $=\sigma_{1}^{j}Z_{j+1}^{(r-2)}\sigma_{2}^{j+r}-\sigma_{2}^{j}Z_{j+1}^{(r-2)}\sigma_{1}^{j+r}.$ (9)
$Z_{j}^{(r)}=\sigma_{3}^{j}...\sigma_{3}^{j+r-1}$ defines a string of $\sigma_{3}$ operators, and $A_{j}^{(1)}(t)=-\sigma_{3}^{j}(t)$. It can be shown Žnidarič and Horvat (2013) that the dynamics of the two point correlation matrix is given by the equation,
@@
Eq. (10) will be the central focus of our analysis and in the next section we show that this equation yields a simple solution via the transfer matrix approach. While we consider a simple case of homogeneous and nearest neighbor couplings, one can carry out the same approach for Hamiltonians with different couplings for odd and even bonds. Also, for other Hermitian Lindblad operators such as the nonlocal dephasing of Ref. Wang et al., 2024 and Ref. Žnidarič, 2024 where the two point correlators still satisfy a closed set of linear equations of the type similar to Eq. (10).
@@ III Transfer Matrix Approach
where $l=x-y$ such that $l=-(N-1),-(N-2),...,-1,0,1,...,N-1$. Using Eq. 12 in Eq. (11) we obtain,
$\frac{d}{dt}g_{l}(t,q)$ $=\mathcal{A}_{ll^{\prime}}g_{l^{\prime}}(t,q)$ (13)
@@
$\mathcal{A}_{ll^{\prime}}=i\frac{\omega(q)}{2}(\delta_{l,l^{\prime}+1}+\delta_{l,l^{\prime}-1})+4\gamma(\delta_{l,0}-1)\delta_{l,l^{\prime}},$
where $\omega(q)=8J\sin(q/2)$.
@@
where $\lambda_{\nu}$ are the eigenvalues of $\mathcal{A}$ and $U$ consists of the corresponding eigenvectors of $\mathcal{A}$. Note that in this solution the dependence on the system size is implicit into the behavior of eigenvectors and the eigenvalues with the system size $L$. Therefore, taking $L\rightarrow\infty$ requires finding the limiting behavior of the eigenvalues and the components of the eigenvectors which in Ref. Eisler, 2011 reduces down to solving a rather complicated Bethe equation for $L\rightarrow\infty$. We will see that the transfer matrix approach avoids the cumbersome calculation of the entire spectrum and the eigenvectors of $\mathcal{A}$ and gives a solution where the dependence on the system size is explicit due to which the limit $L\rightarrow\infty$ can be carried out with ease. It also provides the solution for the full correlation matrix in the limit $L\rightarrow\infty$ where as Ref. Eisler, 2011 considers only the diagonal elements of the correlation matrix in this limit.
We consider Eq. (13) only for $l>0$ as $C_{x,y}(t)$ is Hermitian. Note that this equation simply corresponds to a nearest neighbor tight-binding lattice with the hopping strength of $\omega(q)/2$, and an onsite imaginary chemical potential of $4\gamma$ present at all sites except at $l=0$. Therefore, the standard transfer matrix formalism Lee and Joannopoulos (1981a, b) for tight-binding systems can be applied. However, this formalism is applied to tight-binding systems with boundaries therefore Eq. (13) for $l=0$, which reads
@@
needs to be modified such that it does not contain $g_{-1}(t,q_{n})$. For this purpose, we exploit another symmetry allowed by the equations for the correlators. We note that if initial condition, $C_{xy}(0)$, is such that the even diagonals of the correlation matrix are real and the odd diagonals are imaginary, then this is true for any time (See appendix A). The simplest initial state which guarantees this structure is the initial state where the system is in a product state in $\sigma^{z}$ basis i.e. $\rho(0)$ is a pure state given by $\rho(0)=|\psi_{0}\rangle\langle\psi_{0}|$ where
$|\psi_{0}\rangle=|b_{1}\rangle\otimes|b_{2}\rangle...|b_{L-1}\rangle\otimes|b_{L}\rangle.$ (16)
@@
$\tilde{g}_{l}(s,q_{n})=\mathcal{G}_{ll^{\prime}}(s)g(0,q_{n}),$ (21)
@@
defines Green’s function in the Laplace domain. We will see shortly that for product initial states of the type in Eq. (16), only the first component of the matrix $\mathcal{G}(s)$, $\mathcal{G}_{0,0}(s)$, determines the entire solution and can be expressed as a product of simple $2\times 2$ transfer matrices.
Let us first consider the solution for the diagonal elements of $C(t)$ and consider the remaining elements by the end of this section. The solution for diagonal elements can written in terms of $\tilde{g}_{0}(s,q_{n})$ as,
@@
where $\eta$ is taken such that it is greater than real parts of all the singularities of the function $f(s)$. As the initial condition is diagonal $g_{l}(0,q_{n})=\delta_{l,0}c(q_{n})$, $c(q_{n})$ is the Fourier coefficient of $c_{x}$ i.e.
@@
Using Eq. (26) in Eq. (23) we see that the diagonal elements are determined by the first component $\mathcal{G}_{0,0}(s)$ of the matrix $\mathcal{G}(s)$.
$\mathcal{G}_{0,0}(s)$ can expressed as product of $2\times 2$ matrices using the transfer matrix approach. To that end, we consider the the first column of equations from the identity
@@
namely Eq. (53-55). We show in appendix B that these equations can be used to obtain the following relation,
@@
$\mathcal{G}_{0,0}=\bigg{[}\frac{\tau_{L}}{\sin(zL)}+\frac{1}{\sin(z)}\bigg{[}-\tau_{L-1}+\frac{\tau_{L}\sin(z(L-1))}{\sin(zL)}\bigg{]}\bigg{]}^{-1},$ (30)
where $\tau_{L}=s\sin(zL)-i\omega(q_{n})\sin(z(L-1))$, and $z=\arccos(-iu)$.
Due to the simplicity of the model, we were able to get a simple expression for the required component $\mathcal{G}(s)$ for any finite $L$ and of-course explicit solutions such as in Eq. (30) may not be always possible. Nevertheless, the product structure of Eq. (28) is always possible to obtain as long as there are finitely many non-zero off-diagonal elements in $A$, and the solution is such that even and odd diagonals are purely real and complex, respectively. For example, for models with different hopping on even and odd sites, the transfer matrix will still be $2\times 2$ however the product structure will now contain two different matrices occurring alternately. The dimensions of the transfer matrix is determined by the non-locality of the couplings present and therefore for models with couplings only up to a few neighbors, the dimensions of the transfer matrix will of the same order. For example, the non-local dephasing model with three site Lindblad operators Wang et al. (2024); Žnidarič (2024) (resulting in superdiffusion) defined as $L_{j}=l_{j}^{\dagger}l_{j}$ where,

### v2

@@ Abstract
Solving for the time evolution of a many particle system whose dynamics is governed by Lindblad equation is hard. We extend the use of the transfer matrix approach to a class of Lindblad equations that admit a closed hierarchy of two point correlators. An example that we treat is the XX spin chain, i.e., free fermions, subject to the local on-site dephasing, but can be extended to other Hermitian dissipators, e.g., non-local dephasing. We find a simple expression of the Green’s function in the Laplace domain. The method can be used to get analytical results in the thermodynamic limit, for instance, to get the evolution of the magnetization density and to explicitly see the crossover between ballistic and diffusive behavior, or to show that the correlations between operators at distance $l$ decay with time as $1/t^{\lceil l/2\rceil+1/2}$. It also provides a fast numerical method to determine the evolution of the density with a complexity scaling with the system size more favorably than in previous methods, easily allowing one to study systems with $\sim 10^{6}$ spins.
@@ I Introduction
The transfer matrix formalism has been applied to solve wide variety of problems in physics. A few notable examples include propagation of electromagnetic waves, acoustic waves Yeh (2012); Pendry and MacKinnon (1992); Pendry et al. (1992), waves of quantum particles such as electrons across any scatterers Mello and Kumar (2004); Beenakker (1997); Kirkman and Pendry (1984); Ando (1989); Pendry (1982, 1984); Lee and Joannopoulos (1981a, b), and computing partition functions Izergin et al. (1992); Schultz et al. (1964). This method applies when the solution of the problem can be built iteratively using products of matrices. Therefore, it naturally becomes useful while dealing with problems requiring solutions of linear equations having finitely many non-zero off-diagonals, e.g. tri-diagonal or block tri-diagonal matrices Akaike (1973); Molinari (1997); Reuter and Hill (2012) which describe the Hamiltonian of non-interacting single particle systems. Some simple examples of such systems include the short-ranged quantum tight-binding models or classical harmonic systems. For such systems the transfer matrix formalism allows one to obtain the elements of the single particle Green’s function for the system Lee and Joannopoulos (1981b); Reuter et al. (2011), which determines properties such as the two-point correlation functions, current, conductance etc Dhar and Sen (2006); Dhar and Roy (2006).
@@
Another class of models where such problems occur are quantum mechanical models in presence of dissipation or dephasing due to certain type of Hermitian Lindblad operators Žnidarič (2010); Žunkovič (2014); Barthel and Zhang (2022); Wang et al. (2024); Žnidarič (2011, 2024); Eisler (2011); Temme et al. (2012). In these cases, the dynamics of two point correlation matrix is given by a linear differential equation where the matrix governing the dynamics has finitely many non-zero off-diagonals. One should therefore be able to apply the transfer matrix method in a manner similar to tight-binding lattices to obtain solutions for such equations, and this is the approach taken in this paper. The transfer matrix approach avoids the calculation of the entire spectral decomposition of the matrix governing the dynamics of the correlators and instead expresses the solution as a power of a simple transfer matrix of some small dimension. The latter allows one to take the thermodynamic limit with ease. In this paper, we demonstrate the transfer matrix approach in the simple case of XX spin chain with $L$ spins in the presence of local dephasing and show that this approach leads to a simple solution for the two-point correlation matrix. We express the solution for the correlators in terms of $L^{th}$ power of a $2\times 2$ transfer matrix, and therefore the thermodynamic limit just requires determining its largest eigenvalue and eigenvector. Once the correlation matrix is known, its elements determine physical observables of general interest. For example, the diagonal elements of the correlation matrix directly give the magnetization density and the imaginary parts of the first off-diagonal elements give the magnetization current between different bonds in the chain. Using the transfer matrix approach, we express the Green’s function for the magnetization density as a rather simple function in the Laplace domain whose analytic structure determines the type of the behavior we observe. At long times diffusion arises due to contribution from real poles of the Green’s function while at short times the ballistic behavior comes from the contribution of its branch cut.
@@
XX spin chain with dephasing has been extensively studied in the literature, e.g. Refs. Žnidarič, 2010; Medvedyeva et al., 2016; Teretenkov and Lychkovskiy, 2024; Silva et al., 2023; Cao et al., 2019; Turkeshi and Schiró, 2021; Žnidarič, 2015; Carollo et al., 2017; Bernard and Jin, 2019; Haga et al., 2023, and therefore serves as an ideal simple case that one can consider, and where one can study e.g. a crossover from ballistic magnetization transport at short times to diffusive behavior at long times. One can think of it as the simplest solvable model displaying such a transition. In the continuum limit, a qualitative understanding of the crossover can be described by considering two conserved quantities, say energy and particle density, that are coupled through a scattering. In our spin language we will have the magnetization density $m(x)$ and the magnetization current $j(x)$, with the simplest possible two coupled continuity equations being,
@@
which is nothing but the so called telegrapher’s equation, commonly used in signal analysis for transmission and propagation of electrical signals Metaxas et al. (1983). For small times the diffusive time derivative can be neglected and the equation will behave as a wave equation with velocity $v^{2}$, while for long times the second time derivative is negligible and it goes into a diffusion equation with diffusion constant $D=v^{2}/(4\gamma)$ (see e.g. Ref. Kac, 1974 for a simple approach to solve telegrapher’s equation). Another way to look at it is that it corrects violation of causality in diffusion equation by introducing relaxation time Kadanoff and Martin (1963) (Green’s function for diffusion equation is nonzero everywhere in space already for arbitrarily small times). As we shall see, the equations for the XX chain with dephasing are quite similar, and one can view it as a lattice version of the telegrapher’s equation with a tight-binding dispersion relation, i.e. there are particles with different velocities. What we will get out of our analysis is a behavior that continuously transitions with time from ballistic to diffusive transport, or, in mathematical terms in the continuum limit from a hyperbolic to a parabolic partial differential equation.
This paper is structured as follows: In sec. II, we introduce the model and the matrix equation for the two point correlators that we want to solve. In the next section, sec. III, we adopt the transfer matrix approach to obtain the solution for the correlation matrix at time $t$ for any finite size $L$. In the following section, sec. IV we look at the solution in the thermodynamic limit and discuss the asymptotic behavior of the diagonal and the off-diagonal elements of the correlation matrix. We also present numerical results for the transferred magnetization and its logarithmic derivative (finite-time transport dynamical exponent) and discuss the crossover from the ballistic to diffusive behavior. We conclude in section  V.
@@ II XX-Model with Dephasing
In this section, we introduce the model and setup known equations for the two point correlators that we wish to solve. We consider an $XX$ spin chain with periodic boundary conditions. The Hamiltonian is given by,
@@
Solving the master equation directly requires solving $14^{L}-$ 1 coupled differential equations, and is therefore hopeless for large system sizes. However, certain Eisler (2011); Žunkovič (2014); Barthel and Zhang (2022) Hermitian Lindblad operators with quadratic Hamiltonians allow reducing the exponential complexity in system size to a polynomial complexity. This is achieved by decoupling blocks of observables based on the number of Fermionic operators they contain resulting in a hierarchy of equations. In this hierarchy, lower point correlators serve as source terms for the equations of higher point correlators. For example, the block of two point correlators decouples from the higher point correlators and the two point correlators satisfy a closed set of linear equations Žnidarič (2010, 2011); Žunkovič (2014); Barthel and Zhang (2022); Eisler (2011). These then serve as source terms for third order correlators and so on.
@@
$A_{j}^{(r)}(t)$ $=\sigma_{1}^{j}Z_{j+1}^{(r-2)}\sigma_{1}^{j+r-1}+\sigma_{2}^{j}Z_{j+1}^{(r-2)}\sigma_{2}^{j+r-1},$ (8)
$B_{j}^{(r)}(t)$ $=\sigma_{1}^{j}Z_{j+1}^{(r-2)}\sigma_{2}^{j+r-1}-\sigma_{2}^{j}Z_{j+1}^{(r-2)}\sigma_{1}^{j+r-1}.$ (9)
$Z_{j}^{(r)}=\sigma_{3}^{j}...\sigma_{3}^{j+r-1}$ defines a string of $\sigma_{3}$ operators ($Z_{j}^{(0)}=\mathbbm{1}$), and $A_{j}^{(1)}(t)=-\sigma_{3}^{j}(t)$. It can be shown Žnidarič and Horvat (2013) that the dynamics of the two point correlation matrix is given by the equation,
@@
Eq. (10) will be the central focus of our analysis and in the next section we show that this equation yields a simple solution via the transfer matrix approach. While we consider a simple case of homogeneous and nearest neighbor couplings, one can carry out the same approach for Hamiltonians with different couplings for odd and even bonds and for other Hermitian Lindblad operators such as the nonlocal dephasing of Ref. Wang et al., 2024 and Ref. Žnidarič, 2024 where the two point correlators still satisfy a closed set of linear equations of the type similar to Eq. (10).
@@ III Transfer Matrix Approach
where $l=x-y$ such that $l=-(L-1),-(L-2),...,-1,0,1,...,L-1$. Using Eq. (12) in Eq. (11) we obtain,
$\frac{d}{dt}g_{l}(t,q)$ $=\sum_{l^{\prime}=-(L-1)}^{L-1}\mathcal{A}_{ll^{\prime}}g_{l^{\prime}}(t,q),$ (13)
@@
$\mathcal{A}_{ll^{\prime}}=i\frac{\omega(q)}{2}(\delta_{l,l^{\prime}+1}+\delta_{l,l^{\prime}-1})+4\gamma(\delta_{l,0}-1)\delta_{l,l^{\prime}},$ (14)
with
$\omega(q)=8J\sin(q/2).$ (15)
@@
where $\lambda_{\nu}$ are the eigenvalues of $\mathcal{A}$ and $U$ consists of the corresponding eigenvectors of $\mathcal{A}$. Note that in this solution the dependence on the system size is implicit in the behavior of eigenvectors and the eigenvalues with the system size $L$. Therefore, taking $L\rightarrow\infty$ requires finding the limiting behavior of the eigenvalues and the components of the eigenvectors which in Ref. Eisler, 2011 reduces down to solving a rather complicated transcendental equation for $L\rightarrow\infty$. We will see that the transfer matrix approach avoids the cumbersome calculation of the entire spectrum and the eigenvectors of $\mathcal{A}$ and gives a solution where the dependence on the system size is explicit due to which the limit $L\rightarrow\infty$ can be carried out with ease. It also provides the solution for the full correlation matrix in the limit $L\rightarrow\infty$ whereas Ref. Eisler, 2011 considers only the diagonal elements of the correlation matrix in this limit.
We consider Eq. (13) only for $l>0$ as $C_{x,y}(t)$ is Hermitian. Note that this equation simply corresponds to a nearest neighbor tight-binding lattice with the hopping strength of $\omega(q)/2$, and an onsite imaginary chemical potential of $4\gamma$ present at all sites except at $l=0$. Therefore, the standard transfer matrix formalism Lee and Joannopoulos (1981a, b) for tight-binding systems can be applied. However, this formalism is applied to tight-binding systems with boundaries . Therefore, Eq. (13) for $l=0$, which reads
@@
needs to be modified such that it does not contain $g_{-1}(t,q_{n})$. For this purpose, we exploit another symmetry allowed by the equations for the correlators. We note that if the initial condition $C_{xy}(0)$ is such that the even diagonals of the correlation matrix are real and the odd diagonals are imaginary, then this is true for any time (See appendix A). The simplest initial state which guarantees this structure is the initial state where the system is in a product state in $\sigma^{z}$ basis, i.e. $\rho(0)$ is a pure state given by $\rho(0)=|\psi_{0}\rangle\langle\psi_{0}|$ where
$|\psi_{0}\rangle=|b_{1}\rangle\otimes|b_{2}\rangle...|b_{L-1}\rangle\otimes|b_{L}\rangle,$ (18)
@@
$\tilde{g}_{l}(s,q_{n})=\sum_{l^{\prime}=0}^{L-1}\mathcal{G}_{ll^{\prime}}(s)g_{l^{\prime}}(0,q_{n}),$ (23)
@@
defines Green’s function in the Laplace domain. We will see shortly that for product initial states of the type in Eq. (18), only the first element of the matrix $\mathcal{G}(s)$, $\mathcal{G}_{0,0}(s)$, determines the entire solution and can be expressed in terms of a product of a simple $2\times 2$ transfer matrices.
Let us first consider the solution for the diagonal elements of $C(t)$ and consider the remaining elements at the end of this section. The solution for diagonal elements can be written in terms of $\tilde{g}_{0}(s,q_{n})$ as,
@@
where $\eta$ is taken such that it is greater than the real parts of all the singularities of the function $f(s)$. As the initial condition is diagonal $g_{l}(0,q_{n})=\delta_{l,0}c(q_{n})$, $c(q_{n})$ is the Fourier transform of $c_{x}$ i.e.
@@
Using Eq. (28) in Eq. (25) we see that the diagonal elements are determined by the first element $\mathcal{G}_{0,0}(s)$ of the matrix $\mathcal{G}(s)$.
To express $\mathcal{G}_{0,0}(s)$ in terms of a product of $2\times 2$ matrices, we consider the first column of equations from the identity
@@
namely Eq. (56-58). We show in appendix B that these equations can be used to obtain the following relation,
@@
$\mathcal{G}_{0,0}=\bigg{[}\frac{\tau_{L}}{\sin(\alpha L)}+\frac{1}{\sin(\alpha)}\bigg{[}-\tau_{L-1}+\frac{\tau_{L}\sin(\alpha(L-1))}{\sin(\alpha L)}\bigg{]}\bigg{]}^{-1},$ (32)
where $\tau_{L}=s\sin(\alpha L)-i\omega(q_{n})\sin(\alpha(L-1))$, and $\alpha=\arccos(-iu)$.
III.1 Other Models
Due to the simplicity of the model, we were able to get a simple expression for the required element of the matrix $\mathcal{G}(s)$ for any finite $L$ and of-course explicit solutions such as in Eq. (32) are not always possible. Nevertheless, the product structure of Eq. (30) is always possible to obtain as long as there are finitely many non-zero off-diagonal elements in $A$, and the solution is such that even and odd diagonals are purely real and complex, respectively. For example, for models with different hopping on even and odd sites, the transfer matrix will still be $2\times 2$ however the product structure will now contain two different matrices occurring alternately. The dimensions of the transfer matrix is determined by the non-locality of the couplings present and therefore for models with couplings only up to a few neighbors, the dimensions of the transfer matrix will be of the same order. For example, the non-local dephasing model with three site Lindblad operators Wang et al. (2024); Žnidarič (2024) (resulting in superdiffusion) defined as $L_{j}=l_{j}^{\dagger}l_{j}$ where,